For a certain drilling project, 300 bbls of your prepared mud sample will be required. Determine approximately how many sacks of bentonite Clay (each 100 lbs) will be necessary for the project.

The density of the prepared bentonite + water sample is: 8.55 ppg. Volume of water: 350 cc. Grams of bentonite used: 17.537 grams.
Answer #1

Density= mass/volume

8.55ppg=mass/300bbl

8.55ppg=mass/300×42 gallon

So mass of whole mud and water mixture=8.55×300×42 lbs

=107730lbs

So no of sacks of bentonite clay=107730/100

=1077.3sacks

~1077 sacks